### v7

@@ 5.4.1 Results on FoggyLane
From the Table 2, it is evident that our method achieves state-of-the-art performance on the foggy lane detection dataset FoggyLane, with F1@50, F1@65, F1@75, and F1@85 scores of 95.04, 88.61, 77.32, and 45.02, respectively. These results surpass the best-performing alternative methods by margins of 2.95%, 3.42%, 3.94%, and 3.41%. The mean F1 (mF1) score demonstrates a 3.13% improvement over the strongest competing approach. Notably, the SwinGFFM-t variant achieves an F1@50 score of 92.34, which even exceeds the performance of DLA34-based CLRNet, previously the top-performing method among comparative approaches.
@@ 5.4.2 Results on FoggyCULane
We compared our method with other advanced lane detection methods on the FoggyCULane dataset. As shown in the Table 3, our method consistently achieved the best performance across various scenarios, including Normal, Crowded, shadow, No line, Arrow, Night and Curve. Additionally, our model achieved the highest overall F1 score of 79.85,which achieves a 0.72% improvement over the best alternative method.
@@ 6 Conclusion
Acknowledgments

### v8

@@ 5.4.1 Results on FoggyLane
From the Table 2, it is evident that our method achieves state-of-the-art performance on the foggy lane detection dataset FoggyLane, with F1@50, F1@65, F1@75, and F1@85 scores of 95.04, 88.61, 77.32, and 45.02, respectively. These results surpass the best-performing alternative methods by margins of 2.95%, 4.63%, 5.17%, and 3.41%. The mean F1 (mF1) score demonstrates a 3.32% improvement over the strongest competing approach. Notably, the SwinGFFM-t variant achieves an F1@50 score of 92.34, which even exceeds the performance of DLA34-based CLRNet, previously the top-performing method among comparative approaches.
@@ 5.4.2 Results on FoggyCULane
We compared our method with other advanced lane detection methods on the FoggyCULane dataset. As shown in the Table 3, our method consistently achieved the best performance across various scenarios, including Normal, Crowded, shadow, No line, Arrow and Curve. Additionally, our model achieved the highest overall F1 score of 79.85,which achieves a 0.72% improvement over the best alternative method.
@@ 5.4.3 Results on FoggyTusimple
    Method Venue Backbone F1(%) ACC(%) FP(%) FN(%) SCNN(Pan et al., 2018) AAAI 2018 VGG16 94.64 95.89 5.87 4.78 UFLD(Qin et al., 2020) ECCV 2020 ResNet18 94.57 93.60 3.49 7.59 UFLD(Qin et al., 2020) ECCV 2020 ResNet34 95.05 95.10 4.84 5.08 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet18 95.77 95.78 3.83 3.53 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet34 96.15 95.50 4.59 3.06 LSTR(Liu et al., 2021b) WACV 2021 ResNet34 94.70 95.58 5.74 4.81 RESA(Zheng et al., 2021) AAAI 2021 ResNet34 95.22 95.60 4.63 4.95 RESA(Zheng et al., 2021) AAAI 2021 ResNet50 95.51 95.83 4.53 4.43 CondLane(Liu et al., 2021a) ICCV 2021 ResNet18 95.41 94.81 3.80 5.46 CondLane(Liu et al., 2021a) ICCV 2021 ResNet34 95.65 94.69 3.63 5.13 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet18 96.48 95.64 3.89 3.11 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet34 96.23 95.94 4.34 3.17 FLAMNet(Ran et al., 2023) IEEE TITS 2023 ResNet18 96.52 95.65 3.84 3.09 FLAMNet(Ran et al., 2023) IEEE TITS 2023 ResNet101 96.34 95.98 4.45 2.82 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet18 95.71 93.89 3.53 5.12 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet34 95.61 93.72 3.48 5.38 HWLane(Zhao et al., 2024) IEEE TITS 2024 Res34 95.54 95.91 4.32 4.60 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 ResNet34 96.33 95.91 4.50 2.77 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 DLA34 96.37 94.99 3.21 4.09 Ours SwinGFFM-t 96.63 95.86 2.77 4.04 Ours - SwinGFFM-s 96.95(0.43$\uparrow$) 96.10(0.12$\uparrow$) 2.32(0.89$\downarrow$) 3.83
@@ 6 Conclusion
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgement
@@
Author biography
{biography}
[zrh.jpg]Ronghui Zhang received a B.Sc. (Eng.) from the Department of Automation Science and Electrical Engineering, Hebei University, Baoding, China, in 2003, an M.S. degree in Vehicle Application Engineering from Jilin University, Changchun, China, in 2006, and a Ph.D. (Eng.) in Mechanical & Electrical Engineering from Changchun Institute of Optics, Fine Mechanics and Physics, the Chinese Academy of Sciences, Changchun, China, in 2009. After finishing his post-doctoral research work at INRIA, Paris, France, in February 2011, he is currently an Associate Professor with the Research Center of Intelligent Transportation Systems, School of intelligent systems engineering, Sun Yat-sen University, Shenzhen, 518107, Guangdong, P.R.China. His current research interests include computer vision, intelligent control and ITS.
{biography}
[mayh.jpg]Yuhang Ma received his B.S. degree in Automation at Huazhong University of Science and Technology in 2020. He is currently pursuing the master’s degree in Electronic and Information Engineering at Sun Yat-sen University, Shenzhen, 518107, Guangdong, P.R.China. His current research interests include autonomous driving and computer vision.
{biography}
[ltf.jpg]Tengfei Li received his B.S. degree in Robot Engineering from Harbin Institute of Technology at Weihai in 2024. He is currently pursuing the M.S. degree with the School of Intelligent Systems Engineering, Sun Yat-sen University. His research interests include computer vision and autonomous driving. {biography}[linziyu.pdf]Ziyu Lin is currently pursuing a B.Sc. degree in Traffic Engineering with the School of intelligent systems engineering, Sun Yat-sen University, Guangzhou, China. Her research interests include computer vision, deep learning, and autonomous driving technology, with a particular focus on exploring how these technologies can enhance the efficiency and safety of autonomous driving.
{biography}
[wyy.jpg]Yueying Wu received her B.S. degree in Internet of Things Engineering from Central South University, Changsha, China, in 2020, and an M.S. degree in Transportation Engineering from Sun Yat-sen University, Shenzhen, 518107, Guangdong, P.R.China, in 2023. Her current research interests include autonomous driving and vehicle-road collaboration, deep learning and computer vision. {biography}[cjz.png]Junzhou Chen received his Ph.D. in Computer Science and Engineering from the Chinese University of Hong Kong in 2008, following his M.Eng degree in Software Engineering and B.S. in Computer Science and Applications from Sichuan University in 2005 and 2002, respectively. Between March 2009 and February 2019, he served as a Lecturer and later as an Associate Professor at the School of Information Science and Technology at Southwest Jiaotong University. He is currently an associate professor at the School of Intelligent Systems Engineering at Sun Yat-sen University. His research interests include computer vision, machine learning, intelligent transportation systems, mobile computing and medical image processing.
{biography}
[…]